OpenStudy (anonymous):

Solve the equation for y. 5x-7y=10 A. y= -7/5x - 7/10 B. y= 7/5x + 10/5 C. y= 5/7x - 10/7 D. y= -5x -7

OpenStudy (godlovesme):

Have u tried joey? :)

OpenStudy (anonymous):

Yeah, but math is not my best subject so I have no clue what I'm doing haha. I'm just trying to catch up so I don't fail. :O haha

OpenStudy (godlovesme):

Lol it's okay I'll help ya :) Let me see what u have so far

OpenStudy (igreen):

5x - 7y = 10 First step, subtract 5x to both sides, can you do that?

OpenStudy (anonymous):

Yeah haha.

OpenStudy (godlovesme):

... -_-

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

I got 2 and 5

OpenStudy (anonymous):

?

OpenStudy (igreen):

No..we just subtract it to the other side..we don't combine it with anything.

OpenStudy (igreen):

5x - 7y = 10 5x - 5x - 7y = 10 - 5x

OpenStudy (igreen):

5x - 5x = ?

OpenStudy (anonymous):

I'm majorly confused.

OpenStudy (anonymous):

5-5 would be 0 right?

OpenStudy (anonymous):

You there?